tangents and normals
find the equation of a normal to the curve y=x loge x which is parallel to the line 2x – 2y + 3 = 0 

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

y = xlnx
Slope of normal: -1/y’ = 1
y’ = -1
y' = lnx + 1
y' = lnx + 1 = -1
lnx = -2
x = \frac{1}{e^2}
y = \frac{-2}{e^2}
Equation:

(y + \frac{2}{e^2}) = (x - \frac{1}{e^2})